a thermometer with an arbitrary scale, s, of equal division registers -30°S at the ice point and 90°S at the steam point. Calculate the celsius temperature corresponding to 60°S


1
Expert's answer
2021-02-04T14:39:06-0500

Relationship between C and S scale

"\frac{C}{100}=\frac{S+30}{120}"

"C=\frac{5(S+30)}{6}=75^°" C